How many liters of a 80% acid solution must you mix with 10 liters of a 20% acid solution if you want to end up with a 40% acid solution?
Answer by stanbon(75887)

How many liters of a 80% acid solution must you mix with 10 liters of a 20% acid solution if you want to end up with a 40% acid solution?
----------------------------
Equation:
acid + acid = acid
0.80x + 0.20*10 = 0.40(x+10)
-----
80x + 20*10 = 40x + 40*10
40x = 20*10
x = 5 liters (amt. of 80% solution to be added)
===================================================
cheers,
Stan H.